OpenStudy (anonymous):

What exactly are linear equations, I thought I knew until I found graph paper in my homework book. WTF???

OpenStudy (anonymous):

y=mx+b, of course there are different variations on this. But in general a linear equation has one to two variables. None of those variables have an expoenet attached to them that is larger than 1. All points or pairs of values that makes that equation true lie on a straight line when plotted on a graph

OpenStudy (anonymous):

so is 7(x+4)-5(x+3)+(4-x)=0-a linear equation?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

usually I just solve for x

OpenStudy (anonymous):

yes, you can solve for y and get the y intercept which in this case is : (0,-17). But that is one of many point that you can obtain. In Y=mx+b form this equation is like this: y=x+17. See so, if you set x to 0 you get y=-17 (y intercept). But also we can say that x can be any other number and still that point (x,y) would be located on a straight line. Just like (0,-17). Hence linear equation